OpenStudy (anonymous):

Going to bed soon, but if I could get help with this before that would be great! :) The scale model of a Lamborghini sports car is 1:200. If it requires 12 litres of paint for the real car, how much paint is needed for the scale model in mL? (1 litre = 1000mL) These Scale questions really confuse me! :/

jimthompson5910 (jim_thompson5910):

Let's say we had a cube with each side length of 1 unit  what is the surface area of this cube?

OpenStudy (anonymous):

1 Cubic unit

jimthompson5910 (jim_thompson5910):

that's the volume

jimthompson5910 (jim_thompson5910):

I want to know the total area of each face

OpenStudy (anonymous):

oh sorry.. the area is 1*6? so SA= 6

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now let's say I multiplied every dimension by 200 that's what it means when it says the scale is 1:200 the initial is 1 and the larger is 200 times that

jimthompson5910 (jim_thompson5910):

so let's say we now have this much larger cube with side lengths of 200 

jimthompson5910 (jim_thompson5910):

what is the surface area of that new (much larger) cube?

OpenStudy (anonymous):

200*200 = 40000 40000*6 =240000 SA= 240000 ?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so we've gone from 6 to 240,000

jimthompson5910 (jim_thompson5910):

notice how 240,000/6 = 40,000 and how (200)^2 = 40,000 this is no coincidence

jimthompson5910 (jim_thompson5910):

it turns out that if you know the smaller surface area, you multiply by (200)^2 = 40,000 to get the larger surface area or you can go in reverse: if you know the larger surface area, you can divide by 200^2 = 40,000 to get the smaller surface area

jimthompson5910 (jim_thompson5910):

This works all based on ratios. In general, if you have a cube with side length of 1 and then it jumps to side length of k, then... SA of cube with side length 1: SA = 6*1^2 = 6 SA of cube with side length k: SA = 6*k^2 = 6k^2 Ratio of larger to smaller: (6k^2)/6 = k^2 ---> ratio is k^2 : 1 Ratio of smaller to larger: 6/(6k^2) = 1/(k^2) ---> ratio is 1 : k^2

OpenStudy (anonymous):

Okay! So is 40,000 the SA of the model? And 240 000 the SA of the real car?

jimthompson5910 (jim_thompson5910):

not quite, the 40,000 is based off the example of the cube I drew

jimthompson5910 (jim_thompson5910):

oh nvm

jimthompson5910 (jim_thompson5910):

but 40,000 isn't the answer it helps us get there though

jimthompson5910 (jim_thompson5910):

so we know the entire real car takes 12 liters of paint 12 L = 12*1000 mL 12 L = 12000 mL So it takes 12,000 mL of paint